A random sample of 1400 voters in a particular city found 322 voters who voted yes on proposition 200. Find a 95% confidence interval for the true percent of voters in this city who voted yes on proposition 200.  Express your results to the nearest hundredth of a percent.

Answer:___to___ %
Answer #1

From standard normal tables, we have:
P(-1.96 < Z < 1.96) = 0.95

The sample proportion here is computed here as:
p = x/n = 322/ 1400 = 0.23

Therefore the confidence interval here is computed as:

p \pm z^*\sqrt{\frac{p(1-p)}{n}}

0.23 \pm 1.96*\sqrt{\frac{0.23(1-0.23)}{1400}}

0.23 \pm 0.0220

(0.2080, 0.2520)

This is the required 95% confidence interval here.